Find (g·f)(x) and (f·g)(x).

f(x) = 2x^2+x-7, g(x) = -3x-1

Answer by Boreal(15235)

f(x) = 2x^2+x-7, g(x) = -3x-1
fog=2(-3x-1)^2+(-3x-1)-7
=2(9x^2+6x+1)-3x-8
=18x^2+12x+2-3x-8
=18x^2+9x-6
--------------------------------
gof=-3(2x^2+x-7)-1
=-6x^2-3x+21-1
=-6x^2-3x+20